OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). @ttb123456789 @Nnesha

Nnesha (nnesha):

parallel lines have same slope

Nnesha (nnesha):

firs you have to solve this 4x - 3y = -12 solve this for y

OpenStudy (anonymous):

yes well give me a minute

OpenStudy (anonymous):

y=4/3x+b 4=4/3x*-3+b b=4+4=8 y=4/3x+8

Nnesha (nnesha):

so first subtract 4x both  lside let me know what you get